OpenStudy (anonymous):

solve -2(3y+1)=3(1-6y)-9

OpenStudy (anonymous):

1. Distribute: -6y-2=3-18y-9 2. Combine like terms: 12y-2=-6 3. " ": 12y=-4 4. Divide both sides by 12: y= -4/12 5. simplify: y= -1/3

OpenStudy (anonymous):

Thanks that helped a lot

OpenStudy (anonymous):

glad to help